​1 kg block vibrates in shm with amplitude of 15 cm and period of T = 2 s find the frequency, its maximum speed and its speed

when the displacement is 5cm, its maximum acceleration when the displacement is 5cm.

 


1
Expert's answer
2021-12-09T09:22:45-0500

Frequency:


"f=\frac 1T=0.5\text{ Hz}."

Maximum speed:


"v_\text{max}=2\pi fA=0.47\text{ m/s}."


Speed at x=5 cm:


"x(t)=5=15\cos(2\pi ft),\\
t=0.39\text{ s},\\
v=-v_\text{max}\sin(2\pi ft)=0.44\text{ m/s}."

Acceleration when the displacement is 5 cm:


"a=-(2\pi f)^2A\sin(2\pi ft)=-1.39\text{ m/s}^2."